A cylinder with a piston holds 0.50 m^3 of oxygen at an absolute pressure of 4.0 atm. The piston is pulled outward, increasing the volume of the gas until the pressure drops to 1.0 atm. If the temperature stays constant, what new volume does the gas occupy?
1
Expert's answer
2021-02-15T07:26:40-0500

"V_1 = 0.50 \;m^3 \\

P_1 = 4.0 \;atm \\

P_2 = 1.0 \;atm \\

T = constant \\

P_1V_1 = P_2V_2 \\

V_2 = \frac{P_1 \times V_1}{P_2} \\

= \frac{0.5 \times 4.0}{1.0} \\

= 2.0 \;m^3"

Answer: 2.0 m3